Need help! The universal set contains all from 1-20
Can you please help with answers? Thank you much!
A=(multiples of 3
B=(odd numbers)

Answer by Edwin McCravy(20055)


Need help! The universal set contains all from 1-20
That means the the universal set U is:

U = {1,2,3,4,5,6,7,8,9,10,11,12,13,14,15,16,17,18,19,20}

Can you please help with answers? Thank you much!
A=(multiples of 3

Think of the multiplication facts for 3 that you learned in
3rd grade:

1×3 = 3
2×3 = 6
3×3 = 9
4×3 = 12
5×3 = 15
6×3 = 18
7×3 = 21 <-- we stop with 18 because 21 is more than 20,
and A must contain only numbers that are in U, the universal
set. 

So  A = {3,6,9,12,15,18}

----------------

B=(odd numbers)
The EVEN numbers are the numbers that are or end with 
0,2,4,6, or 8.
The ODD numbers are the numbers that are or end with 
1,3,5,7,or 9.

So for set B, we use only the numbers in U that are or 
end with 1,3,5,7,or 9.

So B = {1,3,5,7,9,11,13,15,17,19}

Edwin